OpenStudy (anonymous):

Solve 81^2+90x=-25 a.) 5/9 b.)-5/9 c.) no real solution d.) plus minus 5/9

OpenStudy (anonymous):

90x = -25 - 81^2 90x = -25 - 6561 Now it should be more easier.

OpenStudy (anonymous):

Now just solve?

OpenStudy (anonymous):

i am going to make a guess my guess is the problem was \[81x^2+90x=-25\] correct me if i am wrong

OpenStudy (anonymous):

in which case you would write \[81x^2+90x+25=0\] and then recognize the "perfect square" on the left and get \[(9x+5)^2=0\] and you should be good from there yes?

OpenStudy (anonymous):

not really :/

OpenStudy (anonymous):

ok then if \[(9x+5)^2=0\] that means that \[9x+5=0\] right?

OpenStudy (anonymous):

and so if \[9x+5=0\] we get \[9x=-5\] and then \[x=-\frac{5}{9}\]

OpenStudy (anonymous):

oh I see! Thanks!

OpenStudy (anonymous):

yw